Find the slope of the line joining points whose coordinates are given
(4,2), (9,5)

Answer by stanbon(75887)

Find the slope of the line joining points whose coordinates are given
(4,2), (9,5)
---------
slope = change in y/change in x = (5-2)/(9-4) = 3/5
------
Cheers,
Stan H.
------------